OpenStudy (anonymous):

If f(x) = 2x − 4 and g(x) = x2 − 4x, find f(4).

OpenStudy (unklerhaukus):

f(x) = 2x - 4 f(4) = 2(4) - 4 = 2×4 - 4 =

OpenStudy (anonymous):

so wat do i do to the problem

OpenStudy (unklerhaukus):

the first step was to plug in 4 for x, then simplify

OpenStudy (anonymous):

so would the answer be 4?

OpenStudy (unklerhaukus):

yes, because f(4) = 2(4) - 4 = 2×4 - 4 = 8 - 4 = 4 good work

OpenStudy (anonymous):

thanks u got yourself a medal and u now have 1140 fans

OpenStudy (unklerhaukus):

\[\huge\color{green}{\ddot\smile}\] is there another part to this question with g(x) in it?

OpenStudy (anonymous):

nope thats it thanks